a uniform steel rod of 1m length is bent in a 90 deg. angle at its mid point.Determine the position of its centre of mass from the corner at bent.

Grade:11

1 Answers

vikas askiitian expert
13 years ago

consider that rod is bent such that half part is on +x axis & other half is on +y axis ...

 

mass of each half is m/2 & length is l/2

 

center of mass of each section will be at mid point of that section so

 

center of mass of rod in x axis will be at (l/4,0)   

center of mass of rod at y axis will be at (0,l/4)

 

now , xcom =[ m1x1 + m2x2]/[m1+m2] =[ (m/2)(l/4) + 0]/[m/2 + m/2]

                  = l/8

    ycom = [m1y1 + m2y2]/[m1+m2] = [(m/2)l/4 + 0]/[m/2+m/2] = l/8

 

rcom = xcom (i) + ycom (j)

        =l/8 i + l/8 j

magnitude = l/4root2                     (  from the corner)